OpenStudy (anonymous):

Use linear approximation to estimate the amount of paint in cubic centimeters needed to apply a coat of paint 0.050000 cm thick to a hemispherical dome with a diameter of 80.000 meters.

OpenStudy (saifoo.khan):

Volume of sphere = 4/3 * π * r^3 dV/dr = 4 * π * r^2 dV = 4 * π * r^2 * dr = 4 * π * 306.25 * .0005 ≈ 1.92 m^3

OpenStudy (anonymous):

would it not be the area we are looking for instead of the volume?

OpenStudy (anonymous):

Cubic centimeters = volume Square centimeters = area

OpenStudy (anonymous):

saifoo.khan that was not the right answer. Any other idea?

OpenStudy (anonymous):

I got it, never mind. It was dV=(3(2/3)*pi*4000^2)*0.05